# Jung Lee Jung Roh Han Cho nocounter]Samsung ResearchSeoulSouth Korea

### v1

@@ Abstract
For text enrollment-based open-vocabulary keyword spotting (KWS), acoustic and text embeddings are typically compared at either the phoneme or utterance level. To facilitate this, we optimize acoustic and text encoders using deep metric learning (DML), enabling direct comparison of multi-modal embeddings in a shared embedding space. However, the inherent heterogeneity between audio and text modalities presents a significant challenge. To address this, we propose Modality Adversarial Learning (MAL), which reduces the domain gap in heterogeneous modality representations. Specifically, we train a modality classifier adversarially to encourage both encoders to generate modality-invariant embeddings. Additionally, we apply DML to achieve phoneme-level alignment between audio and text, and conduct comprehensive comparisons across various DML objectives. Experiments on the Wall Street Journal (WSJ) and LibriPhrase datasets demonstrate the effectiveness of the proposed approach.
@@ 3.2 Utterance-Level Matching
For utterance-level matching, highlighted in the blue area in Fig. 1, we apply pooling operations to $\bm{E}^{\textrm{phn}}_{a}$ and $\bm{E}^{\textrm{phn}}_{t}$ to generate utterance-level embeddings, $\bm{E}^{\textrm{utt}}_{a}\in\mathbb{R}^{d}$ and $\bm{E}^{\textrm{utt}}_{t}\in\mathbb{R}^{d}$. Specifically, channel- and context-dependent statistics pooling (CCSP) [24] is used for $\bm{E}^{\textrm{phn}}_{a}$, while global average pooling is applied to $\bm{E}^{\textrm{phn}}_{t}$. The utterance-level DML loss ($\mathcal{L}_{\textrm{utt}}$) is defined using RP loss. The primary objective of RP Loss is to transfer structural information from TEs to AEs by preserving distance- and angle-wise relationships among TEs. Following [13], we employ all three variants of RP loss: distance-wise RP ($\mathcal{L}_{\textrm{RPL-D}}$), angle-wise RP ($\mathcal{L}_{\textrm{RPL-A}}$), and prototypical RP ($\mathcal{L}_{\textrm{RPL-P}}$) losses.
In [13], the authors demonstrated that incorporating a keyword classification loss ($\mathcal{L}_{\textrm{key}}$) as an auxiliary objective for the acoustic encoder significantly enhances open-vocabulary KWS performance. In this work, we adopt the SphereFace2 framework, originally proposed for face recognition [21] and later adapted for speaker verification [25]. Unlike conventional softmax normalization, which operates under a closed-set assumption, SphereFace2 loss utilizes multiple binary classifiers to train the model in a pair-wise manner, rather than performing multi-class classification. Building on prior studies demonstrating that SphereFace2 outperforms multi-classification-based loss functions, we extend this framework to open-vocabulary KWS and show that SphereFace2 is also effective for this task. To implement this, we convert the PyTorch-based official code provided by [25] [^1] https://github.com/Hunterhuan/sphereface2_speaker_verification into TensorFlow.
@@ 4.5 Performance Comparison on LibriPhrase
We evaluate the performance on the LibriPhrase dataset, which consists of a hard set LPH and an easy set LPE. It is important to note that the models are trained on the King-ASR-066 dataset, not the LibriPhrase training set. As in the previous subsections, to ensure a fair comparison, all methods share the same experimental setup, including datasets and model architectures. The proposed method corresponds to the best model from Table 3, referred to as “ADML+SF2.” We observe that our approach achieves the best performance on the LibriPhrase dataset, consistent with the results on the WSJ dataset.

### v2

@@ Abstract
For text enrollment-based open-vocabulary keyword spotting (KWS), acoustic and text embeddings are typically compared at either the phoneme or utterance level. To facilitate this, we optimize acoustic and text encoders using deep metric learning (DML), enabling direct comparison of multi-modal embeddings in a shared embedding space. However, the inherent heterogeneity between audio and text modalities presents a significant challenge. To address this, we propose Modality Adversarial Learning (MAL), which reduces the domain gap in heterogeneous modality representations. Specifically, we train a modality classifier adversarially to encourage both encoders to generate modality-invariant embeddings. Additionally, we apply DML to achieve phoneme-level alignment between audio and text, and conduct extensive comparisons across various DML objectives. Experiments on the Wall Street Journal (WSJ) and LibriPhrase datasets demonstrate the effectiveness of the proposed approach.
@@ 3.2 Utterance-Level Matching
For utterance-level matching, highlighted in the blue area in Fig. 1, we apply pooling operations to $\bm{E}^{\textrm{phn}}_{a}$ and $\bm{E}^{\textrm{phn}}_{t}$ to generate utterance-level embeddings, $\bm{E}^{\textrm{utt}}_{a}\in\mathbb{R}^{d}$ and $\bm{E}^{\textrm{utt}}_{t}\in\mathbb{R}^{d}$. Specifically, channel- and context-dependent statistics pooling (CCSP) [24] is used for $\bm{E}^{\textrm{phn}}_{a}$, while global average pooling is applied to $\bm{E}^{\textrm{phn}}_{t}$. The utterance-level DML loss ($\mathcal{L}_{\textrm{utt}}$) is defined using RP loss. The primary objective of RP Loss is to transfer structural information from TEs to AEs by preserving distance- and angle-wise relationships among TEs. Building on our previous work [13], we employ all three variants of RP loss: distance-wise RP ($\mathcal{L}_{\textrm{RPL-D}}$), angle-wise RP ($\mathcal{L}_{\textrm{RPL-A}}$), and prototypical RP ($\mathcal{L}_{\textrm{RPL-P}}$) losses.
In [13], we demonstrated that incorporating a keyword classification loss ($\mathcal{L}_{\textrm{key}}$) as an auxiliary objective for the acoustic encoder significantly enhances open-vocabulary KWS performance. In this work, we adopt the SphereFace2 framework, originally proposed for face recognition [21] and later adapted for speaker verification [25]. Unlike conventional softmax normalization, which operates under a closed-set assumption, SphereFace2 loss utilizes multiple binary classifiers to train the model in a pair-wise manner, rather than performing multi-class classification. Building on prior studies demonstrating that SphereFace2 outperforms multi-classification-based loss functions, we extend this framework to open-vocabulary KWS and show that SphereFace2 is also effective for this task. To implement this, we convert the PyTorch-based official code provided by [25] [^1] https://github.com/Hunterhuan/sphereface2_speaker_verification into TensorFlow.
@@ 4.5 Performance Comparison on LibriPhrase
We evaluate the performance on the LibriPhrase dataset, which consists of a hard set LPH and an easy set LPE. To evaluate generalization, all models are trained on King-ASR-066 and tested on LibriPhrase without access to its training set. As in previous subsections, to ensure a fair comparison, all methods share the same experimental setup, including datasets and model architectures. The proposed method corresponds to the best model from Table 3, referred to as “ADML+SF2.” We observe that our approach achieves the best performance on the LibriPhrase dataset, consistent with the results on the WSJ dataset.